OpenStudy (anonymous):

Hey. I'm in Algebra I in 8th grade, and I do really bad. Any tips or anything?

OpenStudy (anonymous):

on what specific topic most important thing is what you do to one side you must do to the other side

OpenStudy (anonymous):

Right now, my class is on percents/decimals, and I can't do that kind of stuff to save my life.

OpenStudy (anonymous):

Percents : say if you are asked to find 15% of 100 - KEY is thata word OF means MULtiPLY =.15 * 100 = 15 let see a prob from your book on that

OpenStudy (anonymous):

Hold on. I'm doing a project right now. I'll be right back and I will explain what I mean.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I don't have any problems in my book that I haven't completed in class... :/

OpenStudy (anonymous):

ok. let me just mention on decimals: if you are given a number and asked to but it in decimal form take the # put it over 100 and simplfy - answer is the # moved over 2 decimal points to the left. note that decimal is at the end of last digit for instance 15% = 15/100 = .15 125% = 125/100 = 1.25